Morrow Corporation had only one job in process during May—Job X32Z—and had no finished goods inventory on May 1. Job X32Z was started in April and finished during May. Data concerning that job appear below:        

Job X32Z
Beginning balance $ 5,100
Charged to the job during May
Direct materials $ 8,200
Direct labor $ 2,100
Manufacturing overhead applied $ 4,100
Units completed 100
Units in process at the end of May 0
Units sold during May 45

In May, overhead was overapplied by $310. The company adjusts its cost of goods sold every month for the amount of the overhead that was underapplied or overapplied.  

Required:  

1. Using the direct method, what is the cost of goods sold for May?

     

   

2. What is the total value of the finished goods inventory at the end of May?

     

3. What is the total value of the work in process inventory at the end of May?
Answer #1

1) cost of goods sold

Beginning balance 5100
Direct material 8200
Direct labor 2100
Manufacturing overhead applied 4100
Total 19500
Unit 100
Unit cost 195

Cost of goods sold = 45*195-310 = 8465

2) Finished goods = 55*195 = 10725

3) Work in process = $0